Activity at the volcano has continued at about the same level as yesterday. 29 rockfall signals, 7 volcano-tectonic earthquakes and 1 hybrid earthquake were recorded by the seismic stations.
The results of COSPEC measurements completed today indicated a decrease in sulphur dioxide flux from yesterday to 1600 tonnes per day.
The results of occupation of GPS stations yesterday indicate continued small changes at the Hermitage site only.
This morning's rain has helped to keep ash levels low in all areas measured.
Residents should continue listening to Radio Montserrat in case there are more sudden developments in the state of the volcano.